Photochronograph Corporation (PC) manufactures time series photographic equipment. It is currently at its target debt-equity ratio of 0.75. Its considering building a new $54 million manufacturing facility. This new plant is expected to generate aftertax cash flows of $6.6 million in perpetuity. The company raises all equity from outside financing. There are three financing options 1. A new issue of common stock: The flotation costs of the new common stock would be 8.4 percent of the amount raised. The required return on the companys new equity is 15 percent 2. A new issue of 20-year bonds: The flotation costs of the new bonds would be 3.0 percent of the proceeds. If the company issues these new bonds at an annual coupon rate of 6.0 percent, they will sell at par 3. Increased use of accounts payable financing: Because this financing is part of the companys ongoing daily business, it has no flotation costs, and the company assigns it a cost that is the same as the overall firm WACC. Management has a target ratio of accounts payable to long-term debt of 0.10 (Assume there is no difference between the pretax and aftertax accounts payable cost.) What is the NPV of the new plant? Assume that PC has a 38 percent tax rate. (Enter your answer in dollars, not millions of dollars, i.e. 1,234,567. Do not round intermediate calculations and round your final answer to the nearest whole dollar amount.] NPV
Answer #1

SEE SCREENSHOT. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE. THANK YOU.

Home nert Page Layout Formulas Data Review View dd-Ins Cut Copy Σ AutoSum . illSo & Find & 2 ClearFe Select 11.A. A.--- 9- ー E ゴText ずWrap Text General Paste チFormat Painter B l U. ㄧ˙ タ.Δ. -=- 還便困Merge & Center. $,%, ,,o Conditional Format Cell Insert Delete Format as Table Styles Styles Clipboard FB178 ER Font Alignment Number Cells Edting ES ET EU EX EY EZ FA FB FC FD 160 161 162 163 164 165 166 167 168 169 170 Accounts payable to long term debt ratio Accounts payable weight- 0.1 0.090909 W(LTD) 0.909091 W(A/P) ng term debt weight- Debt equity ratio DEBT WEIGHT EQUITY WEIGHT = 0.75/1.75 1/1.75 0.75 0.428571 Wd 0.571429 We Account payable has same cost as that of WACC WACCWe*ke+Wd (W(LTD)*kd(1-t) W(A/P) WACC (1-t)) WACC- 0.571429(15%) +0.428571(0.909091(6%)(10.38) 40,090909#WACC) WACC= 8.571435% + 1.44934996 +0.038961(WACC) 0.961039 (WACC) = 10.02078434% WACC . 172 173 174 175 176 10.4270319 % (1/1.75)*(8.4%) + (0.75/1.75)*(0.909091*(3%)+ 0.090909(0)) 5.968831% FLOTATION COSTS- 178 EV PE EPS DIV Circular References: AL125 PB ROPI EVA MVA LEVERAGED BUY- S AP CAP GAIN 福 トト1 ABNORMAL . APV ROSS debe k MARGIN MONEY ma gral tax -price quote di tess BRIC HAN Ready 130% 21-01-2019Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー! 曰 Wrap Text ta copy. в 1 프 . Ej-., Δ. 遑锂函Merge & Center. $, % , Paste 弼,8 Conditional Format CeInsert Delete Format Formatting, as Table w styles. ▼ ㆆ ▼ Sort &Find & 2 ClearFe Select Edting Format Painter Clipboard FB193 ER Font Alignment Number Cells ES ET EU EV EW EX EY EZ FA FB FC FD 175 176 (1/1.75)*(8.4%) + (0.75/1.75)*(0.909091(3%)+ 0.090909*(0)) 5.968831% FLOTATION COSTS = 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 il 1 r THEREFORE AMOUNT TO BE RAISED 54000000/(1-5.968831%) 57427766.49 NPV = NPV- NPV- PV OF PERPETUAL CASH FLOWS- AMOUNT RAISED (6600000/0.104270319)-57428480.27 5869249 (no intermediate rounding is done) eEVA MVA LEVERAGED BUY SWAP distress BRIGHAN CAP GAIN 1 EV PE EPS DİV ABNORMAL Circular References: AL125 APV ROSS deberk MARGIN MONEY al tax . PB ROPİ Ready 福 130% 17:22 21-01-2019